OpenStudy (anonymous):

how would i work them with numbers can you give me some examples? this is the problem 3/5+5/6

OpenStudy (anonymous):

3/5+5/6, get common denominators by multipying the left side by 6/6 and the right side by 5/5. 3/5(6/6)+5/6(5/5) = (3*6)/(5*6)+(5*5)/(6*5) = 18/30+25/30 = 43/30

OpenStudy (shayaan_mustafa):

Hi asia101! :) still needs help friend?

OpenStudy (anonymous):

yes on a different subject